Two forces equal in magnitude and opposite in dire
Answer #1

(a)

Net torque acting on the rod is,

tow = 3*F1 - (3+L) * F2

5.2= 3*7.9 - (3+L) * 7.9


Solve for L

L= 0.658 m

----------------------------

(b)

Net torque is clockwise.

---------

(c)

Net torque acting on the rod is,

tow =L*F1

5.2 = L * 7.9

L = 0.658 m